OpenStudy (anonymous):

can you please tell me the expanded form of x^4/x^2

OpenStudy (kainui):

Expanded form or simplified form?

OpenStudy (anonymous):

expanded

OpenStudy (anonymous):

and simplified

OpenStudy (anonymous):

please help me

Parth (parthkohli):

This one?\[{x \times x \times x \times x \over x \times x}\]\[\implies {\cancel{x \times x} \times x \times x \over \cancel{x \times x}}\]

OpenStudy (anonymous):

um i guess so

OpenStudy (anonymous):

so the simplified form is x^2

Parth (parthkohli):

Yes :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

but I don't know the expanded form

Parth (parthkohli):

as I said, this is the expanded form\[{x \times x \times x\times x \over x \times x }\]

OpenStudy (anonymous):

oh I'm sorry I didn't know that